OpenStudy (anonymous):

Find vectors perpendicular to both of the given vectors. (a) ---> (3i-2k) and 2k

OpenStudy (anonymous):

do the same cross product: your vectors are (3,0,-2) and (0,0,2)

OpenStudy (anonymous):

-6j

OpenStudy (anonymous):

Oh, Ok.

OpenStudy (anonymous):

It says j is also a possiblity in my textbook?

OpenStudy (anonymous):

ya, any j actualy

OpenStudy (anonymous):

It just says, -6j or j are the answers.. how is j?

OpenStudy (anonymous):

sry.. Only -6j this is what you get when cros multiply your vectors.

OpenStudy (anonymous):

But it says j is also an answer...

OpenStudy (anonymous):

(3,0,-2) and (0,0,2) (0,a,0).(3,0,-2) = 0 + 0 + 0 =0 (0,a,0).(0,0,2) = 0 + 0 + 0 = 0 So actualy any (0,a,0) vector will do. I said -6j, since that's the result of cross multiplication, but this vector is equal -6(0,1,0) no? so it's paralel .

OpenStudy (anonymous):

any paralel vector to -6j will be also a solution. Ok?

OpenStudy (anonymous):

hmmm. OK

OpenStudy (anonymous):

if you dont get it, tell me i try to explain better

OpenStudy (anonymous):

I understand. Don't worry :)

OpenStudy (anonymous):

kk